OpenStudy (anonymous):

Ok i need help with a problem. It says Camilla makes and sells jewelry.She has 8160 silver beads and 2880 black beads.Each necklace will contain 85 silver beads and 30 black beads.How many necklaces can she make?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Thats the answer?How did u get it?

OpenStudy (anonymous):

i made it up how on earth can you answer this?

OpenStudy (anonymous):

Ugh i really need help.

OpenStudy (anonymous):

post the entire question

OpenStudy (fibonaccichick666):

there isn't enough info. You are missing part of the question

OpenStudy (anonymous):

Thats the question

OpenStudy (anonymous):

Oh each necklace contains 85 beads

OpenStudy (anonymous):

then i will stick with 12

OpenStudy (fibonaccichick666):

that would be very necessary

OpenStudy (anonymous):

Im a lil bit dislexic

OpenStudy (anonymous):

a minor detail

OpenStudy (anonymous):

add the beads, divide by 85

OpenStudy (anonymous):

ooh now we have a question !!

OpenStudy (anonymous):

Each necklace contains 85 silver beads and 30 clack beads

OpenStudy (anonymous):

how many times does 85 go in to 8160 ?

OpenStudy (fibonaccichick666):

even more things we need

OpenStudy (anonymous):

85 goes in 8160 96 times

OpenStudy (anonymous):

lol those hits just keep on coming what is \(8160\div 85\) ? use a calculator , round if necessary

OpenStudy (anonymous):

k how many times does 30 go in to 2880?

OpenStudy (anonymous):

hold the phone

OpenStudy (anonymous):

96.So 96 necklaces?

OpenStudy (anonymous):

did you compute \[2280\div 30\]?

OpenStudy (anonymous):

yes it says it equals 96

OpenStudy (anonymous):

damn maybe i am putting it in wrong

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=2280%2F30

OpenStudy (anonymous):

my mistake it equals 76

OpenStudy (anonymous):

oh whew! that is what i got pick the smaller of the two numbers i.e. 76

OpenStudy (anonymous):

ok so that is the answer?

OpenStudy (fibonaccichick666):

just so you know for when you submit assignments @Jassy1002 , make sure you show every minute detail of your work so that if you make a numerical mistake you still get some credit

OpenStudy (anonymous):

what i just said

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

she can make 76 necklaces, then she will run out of black beads and have to go to the bead store for more

OpenStudy (anonymous):

the one that smells like incense and sandalwood

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

you got more or is that all?

OpenStudy (anonymous):

I got more.

OpenStudy (anonymous):

It says how is the quotient 80,000 divided by 2000 different from the quotient 80,000 divided by 200 or divided by 20 diffrent...Its late and i can barely think

OpenStudy (anonymous):

\[\frac{80,000}{2,000}=\frac{80\cancel{000}}{2\cancel{000}}=\frac{80}{2}=40\] you cancel three of the zeros before dividing with \[\frac{80,000}{200}\] you cancel only 2 zeros

OpenStudy (anonymous):

lol it must be late

OpenStudy (anonymous):

So the diffrence is the amount of the zeros???

OpenStudy (anonymous):

yes of course

OpenStudy (anonymous):

So that is all?

OpenStudy (anonymous):

\(2,000\) has three zeros \(200\) has \(2\) and \(20\) has one

OpenStudy (anonymous):

So how is the answer different and the amount of zeros is what is different?

OpenStudy (anonymous):

so when you cancel the zeros (which is really cancelling tens) you cancel either 3, 2, or 1 of them

OpenStudy (anonymous):

not sure what you are asking me \[\frac{80,000}{2000}=\frac{80}{2}=40\] whereas \[\frac{80,000}{200}=\frac{800}{2}=400\] and \[\frac{80,000}{20}=\frac{8000}{2}=4000\]

OpenStudy (anonymous):

Then here is a problem saying Explain how you get from the square to the square root and what it means.

OpenStudy (anonymous):

What is the diffrence between the answers ?

OpenStudy (anonymous):

you gotta write the whole thing does it really just say "Explain how you get from the square to the square root and what it means."

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

well if you have a square, like say \(5^2=25\) is a square then \(\sqrt{25}=\sqrt{5^2}=5\)

OpenStudy (anonymous):

How do you do that symbol?!

OpenStudy (anonymous):

\[\sqrt{\diamondsuit}\]? i am using latex

OpenStudy (anonymous):

But how do i get from a square to a square root?

OpenStudy (anonymous):

if \(a^2\) is the square of \(a\) then \(\sqrt{a^2}=a\)

OpenStudy (anonymous):

for example the square root of \(4^2\) is \(4\)

OpenStudy (fibonaccichick666):

they are inverses jassy, just like 1+(-1)=0 \(\frac{1}{2}*\frac{2}{1}=1\)

OpenStudy (anonymous):

So confusing.

OpenStudy (anonymous):

ok lets unconfuse you pick a number, i don't care what you say (don't make it too big)

OpenStudy (anonymous):

2

OpenStudy (anonymous):

k that is a nice small number now what is \(2^2=2\times 2\) ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

yup and now what is \(\sqrt{4}\) ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

that is not that confusing then is it?

OpenStudy (anonymous):

But i cant use examples i have to explain how i got from a square to a squared root.

OpenStudy (anonymous):

take a number, square it, get another number take the square root of the number you squared, you get your original number back

OpenStudy (anonymous):

the square root of a square, is the number being squared like in your example you squared 2, got 4, take the square root of 4, get 2

OpenStudy (anonymous):

i think you are thinking too hard, making more of this than there is there is not much to write just try to make it plain english

OpenStudy (anonymous):

this some sort of on line class or something ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

oh you got to go turn this in at school then i guess

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

then write something similar to what i wrote and why not include an example? should be allowed if you do in with symbols you can say \(\sqrt{a^2}=a\)

OpenStudy (anonymous):

So what you r saying is that i answer the exponent take that number find the square root of it and that is how i get from a square to a square root?

OpenStudy (anonymous):

Yes no maybe?

OpenStudy (anonymous):

"Explain how you get from the square to the square root and what it means." if the number is a square, like the square of 5 is 25 and the square of \(a\) is \\(a^2\) then the square root of the square is the number being squared, like \(\sqrt{5^2}=5\) and \(\sqrt{a^2}=a\)

OpenStudy (anonymous):

So that is the answer??

OpenStudy (anonymous):

"the exponent take that number find the square root of it " is not something i really understand

OpenStudy (anonymous):

that is an answer, yes it is my best answer at the moment

OpenStudy (anonymous):

seems like a pretty annoying math class what class is it?

OpenStudy (anonymous):

Its 12 AM i need to finish this cause i gotta be up at 5:30 am

OpenStudy (anonymous):

how many more you got?

OpenStudy (anonymous):

Um Adv math or 7th grade math..I dont remember

OpenStudy (anonymous):

you live on a farm? got cows to milk?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

5:30??!!

OpenStudy (anonymous):

got another? lets knock it out so you can get a couple hours sleep

OpenStudy (anonymous):

It takes me an hour to get ready and bout an hour to get to school and i have to be in homeroom by 7:35 am

OpenStudy (anonymous):

Um i think i can do most of these by myself..

OpenStudy (anonymous):

I have like 21 more questions tho

OpenStudy (anonymous):

k good luck

OpenStudy (anonymous):

Wait!

OpenStudy (anonymous):

I need help understanding these instructions.

OpenStudy (anonymous):

k i will wait

OpenStudy (anonymous):

It says Estimate each quotient by rounding the dividend and the divisor to the largest place value.Then estimate each quotient by rounding the dividend and divisor to the nearest 10.

OpenStudy (anonymous):

ok what quotient?